# A priori error analysis of the proximal Galerkin method

### v2

@@ 3.1. Compatible subspaces
The density of $W$ in $Q^{\prime}$ and the closed range theorem can be used to show that 41 implies $B^{\prime}\colon Q^{\prime}\to V^{\prime}$ is bounded from below. Likewise, 40 ensures that $B^{\prime}$ remains bounded from below (uniformly in $h$) after discretization.
To prove 40, it suffices to exhibit a continuous so-called Fortin operator $\Pi_{h}\colon V\rightarrow V_{h}$ satisfying $\|\Pi_{h}v\|_{V}\lesssim\|v\|_{V}$ for all $v\in V$ and
@@
see, e.g., [25, Lemma 26.9] and [7, Section 5.4.3]. In what follows, we let $\|\Pi_{h}\|$ denote the operator norm of $\Pi_{h}$.
@@
The associated LBB condition 41 holds with equality for $\beta=1$, which immediately follows from the definition of the $H^{-1}(\Omega)$ norm. In Section 4, we provide two examples of compatible subspaces $V_{h}\times W_{h}$ satisfying 40 with their corresponding Fortin operators.
@@
The following inf-sup condition holds with a constant $\beta>0$ [14, Proposition 7.2 and Remark 7.2]:
@@
In Section 5, we provide an example of $V_{h}\times W_{h}$ satisfying 40, and we construct the corresponding Fortin operator.
@@
with $Q=L^{2}(\Omega)$. For the Legendre function $\mathcal{R}$, one can choose the Fermi–Dirac entropy given in Example 2.2 with $\underline{u}=-1$ and $\overline{u}=1$, although other convenient choices are also appropriate; cf. [28]. The continuous LBB condition 41 holds thanks to the surjectivity of the divergence operator from $H(\mathrm{div},\Omega)$ to $L^{2}(\Omega)$; see, e.g., [25, Lemma 51.2]. To ensure that 40 holds, a natural choice of subspace would be the $H(\mathrm{div})$-conforming Raviart–Thomas space for $V_{h}$ and the broken polynomial space of the same order for $W_{h}$; see, e.g., [25, Lemma 51.10].
@@ 3.2. The PG method is well-defined
which implies $\hat{u}_{h}=u_{h}$. Equation 10a then implies $b(v_{h},\hat{\psi}_{h}-\psi_{h})=0$ for all $v_{h}\in V_{h}$ and the assumed compatibility assumptions on $V_{h}$ and $W_{h}$ yield $\hat{\psi}_{h}=\psi_{h}$. Indeed, a direct consequence of 40 is that we can estimate
@@
Substituting $v=v(w)\in V_{h}$ into 49, we obtain
@@
Now, employing formula 52 we bound $\|v(w)\|_{V}$ from below:
@@
where, in the first step, we used that the isomorphism $A^{-1}:V_{h}^{\prime}\to V_{h}$ is injective with closed range. In the last estimate, we used 48. Together with the fact that we can lower bound $(\mathcal{R}^{*}(w),1)_{\Omega_{d}}$ by an affine linear function and the equivalence of norms in finite dimensional spaces, we deduce that $J(w)\to\infty$ as $\|w\|_{Q^{\prime}}\to\infty$. The map $w\mapsto v(w)$ is continuous as the first term is affine linear and $\mathcal{R}^{*}$ is continuous, hence we can guarantee the existence of a maximizer $w^{*}$ for $J$.
We now show that the pair $(v(w^{*}),w^{*})$ is the sought-after saddle point of $\mathcal{L}$. This follows from standard arguments; see e.g., [16, Exercise 7.16-4]. We provide some details for completeness. First observe that since $\mathcal{L}(v,\cdot)$ is concave, we have that
@@
where we used that $\mathcal{L}(v(w^{*}),w^{*})\leq\mathcal{L}(v(\delta_{\theta}),w^{*})$, see 50. Then, we obtain that for any $\theta>0$, $\mathcal{L}(v(\delta_{\theta}),w)\leq\mathcal{L}(v(w^{*}),w^{*})$. With the continuity of the map $w\mapsto v(w)$ and of $\mathcal{L}(\cdot,w)$, we conclude that
@@ 3.4. Best approximation error estimates
Thus, upon dividing 63 by $\sum_{k=1}^{\ell}\alpha_{k}$ then adding and subtracting $\langle E^{\prime}(u),\Pi_{h}u-u\rangle$, we obtain
@@
At this stage, we select $u=u^{*}$ and note that for any $v\in K$, 59 and 9 provide the bound
@@
Thus, adding $\langle E^{\prime}(u^{\ast}),v-u_{h}^{\ell}\rangle$ to both sides of 66 and recalling that $Bu^{*}=o^{*}$, we arrive at
@@
We first estimate $\|B^{\prime}(\lambda^{*}-\lambda_{h}^{\ell})\|_{V^{\prime}}$. For any $v\in V$, we use the Fortin operator 42 and equation 10a to write
@@
Since the map $B:V\rightarrow Q$ is surjective, it follows from the closed range Theorem, see e.g. [22, Lemma A.40], that there exists a constant $\beta>0$ with
@@
Combining 69 with 70 finishes the proof. ∎
@@
(71) $\|Bu^{*}-o_{h}^{\ell}\|_{L^{2}(\Omega_{d})}\leq\|B(u^{*}-u_{h}^{\ell})\|_{L^{2}(\Omega_{d})}+\|(I-P_{W_{h}})(Bu_{h}^{\ell}-o_{h}^{\ell})\|_{L^{2}(\Omega_{d})},$
where $P_{W_{h}}$ is the $L^{2}$-projection operator onto $W_{h}$ 75.
@@
This is a direct consequence of the triangle inequality and the observation that $P_{W_{h}}Bu_{h}^{\ell}=P_{W_{h}}o_{h}^{\ell}$, which follows from 10b. ∎
@@ 3.5. Convergence rates and mesh-independence
We now derive abstract error convergence rates from Theorem 3.3 and Lemmas 3.4 and 3.5. For simplicity, we focus on the following spaces $V\subset H^{1}(\Omega)$ and $V\subset H^{1}(\Omega;\mathbb{R}^{n})$. Hereafter, we do not explicitly differentiate between scalar and vector-valued spaces, using the notation for the scalar space while noting that the same results hold, mutatis mutandis, for vector-valued spaces. To derive error rates, we require the construction of an enriching map that allows us to control the last term in (58). In turn, we make the following general assumption, which is critical to defining this map. First, we motivate this assumption with a remark.
@@
(72) $u_{h}^{\ell}\in K_{h}=\left\{v_{h}\in V_{h}\mid\int_{K}(v_{h}-\phi)>0\right\}.$
As expected from the analysis of nonconforming discretizations of obstacle problems [5, Section 5.2.1], one then requires an operator $\mathcal{E}_{h}:K_{h}\rightarrow K$ with suitable approximation properties. See Section 4.3 for an explicit construction of the enriching map for $K_{h}$ defined in 72.
@@
There exists a continuous map $\mathcal{E}_{h}:V\rightarrow V$ with the property:
(73) $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{E}_{h}:K\cup K_{h}\rightarrow K$},
where
(74) $K_{h}=\{u_{h}\in V_{h}\mid P_{W_{h}}(Bu_{h}-\nabla\mathcal{R}^{*}(\psi))=0\text{ for some }\psi\in L^{\infty}(\Omega_{d};\mathbb{R}^{m})\},$
and $P_{W_{h}}$ is the $L^{2}(\Omega_{d};\mathbb{R}^{m})$-projection operator onto $W_{h}$, defined by
(75) $(P_{W_{h}}o,w_{h})_{\Omega_{d}}=(o,w_{h})_{\Omega_{d}}~\text{ for all }o\in L^{2}(\Omega_{d};\mathbb{R}^{m}),~w_{h}\in W_{h}.$
The set $K_{h}$ contains functions $u_{h}\in V_{h}$ such that $Bu_{h}\in W_{h}$ is the $L^{2}$-projection of a constrained observable $o\in\mathcal{O}$. It is constructed to ensure that $u_{h}^{k}\in K_{h}$ for all iterations $k$; see 10b. To derive error rates, we require the following quasi-interpolation assumption on the stability and approximation of the Fortin and enriching operators. The decomposition of the enriching map given in 77 reflects the constructions used in Sections 4 and 5, below.
Assumption 3.
Assume that the Fortin operator satisfying 42 is stable in the sense that
@@
for all $v\in V$. Further, assume that the enriching map $\mathcal{E}_{h}$ of Assumption 2 is affine linear:
@@
Since $E^{\prime}(u^{*})\in H^{r-1}(\Omega)$, we further bound the last two terms in 58 by $\|E^{\prime}(u^{*})\|_{H^{r-1}(\Omega)}\|\Pi_{h}u^{*}-u^{*}\|_{H^{1-r}(\Omega)}$ and by $\|E^{\prime}(u^{*})\|_{H^{r-1}(\Omega)}\|v-u_{h}^{\ell}\|_{H^{1-r}(\Omega)}$ respectively. Thus, we obtain
@@
To handle the $\|u^{*}-\Pi_{h}u^{*}\|_{H^{1-r}(\Omega)}$ term in 80, we first note that from space interpolation between $L^{2}(\Omega)$ and $H^{1}(\Omega)$ (see, e.g., [8, Chapter 14, Proposition 14.1.5]) and 78a, we obtain that for any $w\in H^{1+s}(\Omega)$,
@@
The estimate on $\|\lambda^{*}-\lambda_{h}^{\ell}\|_{H^{-1}(\Omega)}$ follows from Lemma 3.4. In particular, the first term in 68 is bounded by:
@@ 3.6. Discussion of results
  • First, the PG iterates $u^{\ell}_{h}$ converge to the true solution $u^{\ast}$ as $\ell\to\infty$ and $h\to 0$, even with bounded proximity parameters, $\alpha_{k}\leq\textrm{const.}$ This stands in contrast to penalty [30, Chapter 1.7] and interior point methods [38, Chapter 19] — and even augmented Lagrangian methods in infinite-dimensional spaces [35, 3] — which all require taking a relaxation parameter to a singular limit.
  • Second, the number of iterations is asymptotically independent of the mesh size. More specifically, Corollary 1 shows that a user can guarantee convergence to any desired accuracy by independently selecting the mesh size $h$ and iteration count $\ell$. Such mesh-independence is an important feature [45] seen in, e.g., the primal-dual active set method [32], but for a more restrictive class of problems than considered here [34]. Moreover, note that Corollary 1 is a global result, holding for any uniformly bounded initial guess $\psi_{h}^{0}$. In contrast, the celebrated mesh-independence of Newton’s method [1, 50] is only a local property, as it requires an accurate initial guess on a sufficiently fine mesh.
  • Finally, PG can be applied without modification to problems with low-regularity solutions or multipliers. Indeed, Theorem 3.6 provides sufficient conditions on $\Pi_{h}$ and $\mathcal{E}_{h}$ to obtain optimal convergence rates in $h$ depending naturally on the solution $u^{\ast}$ and Fréchet Derivative $E^{\prime}(u^{\ast})$. The following two sections show that these abstract conditions are checkable in practice. It is well-known that penalty methods can also be applied without modification in the low-regularity setting [33]. However, to the best of our knowledge, the existing theory for discretized penalty methods requires $u^{*}\in H^{1+s}(\Omega)$ and $s\in(1/2,1]$ when applied to the obstacle and Signorini problems. See [44, 31] for the obstacle problem and [14, Section 6] for the Signorini problem. Notably, the PG results hold for any $s\in(0,1]$.
@@ 4. Application I: The obstacle problem
and the obstacle $\phi\in H^{1}(\Omega)\cap L^{\infty}(\Omega)$ is fixed. For simplicity, we assume that $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi|_{\partial\Omega}=-\delta\leq 0$} for some non–negative constant $\delta$. It is well-known that there exists a unique solution to 87 [16, Theorem 6.9-1], here denoted by $u^{\ast}$.
@@
Assumption 1 holds for this choice of Legendre function [36, Theorem 4.1]; see also [36, Proposition A.8]. Alternative choices for $\mathcal{R}$ are also admissible, without affecting the proceeding error analysis.

### v3

@@ 3.1. Compatible subspaces
The density of $W$ in $Q^{\prime}$ and the closed range theorem can be used to show that 42 implies $B^{\prime}\colon Q^{\prime}\to V^{\prime}$ is bounded from below. Likewise, 41 ensures that $B^{\prime}$ remains bounded from below (uniformly in $h$) after discretization.
To prove 41, it suffices to exhibit a continuous so-called Fortin operator $\Pi_{h}\colon V\rightarrow V_{h}$ satisfying $\|\Pi_{h}v\|_{V}\lesssim\|v\|_{V}$ for all $v\in V$ and
@@
see, e.g., [27, Lemma 26.9] and [8, Section 5.4.3]. In what follows, we let $\|\Pi_{h}\|$ denote the operator norm of $\Pi_{h}$.
@@
The associated LBB condition 42 holds with equality for $\beta=1$, which immediately follows from the definition of the $H^{-1}(\Omega)$ norm. In Section 4, we provide two examples of compatible subspaces $V_{h}\times W_{h}$ satisfying 41 with their corresponding Fortin operators.
@@
The following inf-sup condition holds with a constant $\beta>0$ [16, Proposition 7.2 and Remark 7.2]:
@@
In Section 5, we provide an example of $V_{h}\times W_{h}$ satisfying 41, and we construct the corresponding Fortin operator.
@@
with $Q=L^{2}(\Omega)$. For the Legendre function $\mathcal{R}$, one can choose the Fermi–Dirac entropy given in Example 2.2 with $\underline{u}=-1$ and $\overline{u}=1$, although other convenient choices are also appropriate; cf. [30]. The continuous LBB condition 42 holds thanks to the surjectivity of the divergence operator from $H(\mathrm{div},\Omega)$ to $L^{2}(\Omega)$; see, e.g., [27, Lemma 51.2]. To ensure that 41 holds, a natural choice of subspace would be the $H(\mathrm{div})$-conforming Raviart–Thomas space for $V_{h}$ and the broken polynomial space of the same order for $W_{h}$; see, e.g., [27, Lemma 51.10].
@@ 3.2. The PG method is well-defined
which implies $\hat{u}_{h}=u_{h}$. Equation 10a then implies $b(v_{h},\hat{\psi}_{h}-\psi_{h})=0$ for all $v_{h}\in V_{h}$ and the assumed compatibility assumptions on $V_{h}$ and $W_{h}$ yield $\hat{\psi}_{h}=\psi_{h}$. Indeed, a direct consequence of 41 is that we can estimate
@@
Substituting $v=v(w)\in V_{h}$ into 50, we obtain
@@
Now, employing formula 53 we bound $\|v(w)\|_{V}$ from below:
@@
where, in the first step, we used that the isomorphism $A^{-1}:V_{h}^{\prime}\to V_{h}$ is injective with closed range. In the last estimate, we used 49. Together with the fact that we can lower bound $(\mathcal{R}^{*}(w),1)_{\Omega_{d}}$ by an affine linear function and the equivalence of norms in finite dimensional spaces, we deduce that $J(w)\to\infty$ as $\|w\|_{Q^{\prime}}\to\infty$. The map $w\mapsto v(w)$ is continuous as the first term is affine linear and $\mathcal{R}^{*}$ is continuous, hence we can guarantee the existence of a maximizer $w^{*}$ for $J$.
We now show that the pair $(v(w^{*}),w^{*})$ is the sought-after saddle point of $\mathcal{L}$. This follows from standard arguments; see e.g., [18, Exercise 7.16-4]. We provide some details for completeness. First observe that since $\mathcal{L}(v,\cdot)$ is concave, we have that
@@
where we used that $\mathcal{L}(v(w^{*}),w^{*})\leq\mathcal{L}(v(\delta_{\theta}),w^{*})$, see 51. Then, we obtain that for any $\theta>0$, $\mathcal{L}(v(\delta_{\theta}),w)\leq\mathcal{L}(v(w^{*}),w^{*})$. With the continuity of the map $w\mapsto v(w)$ and of $\mathcal{L}(\cdot,w)$, we conclude that
@@ 3.4. Best approximation error estimates
Thus, upon dividing 64 by $\sum_{k=1}^{\ell}\alpha_{k}$ then adding and subtracting $\langle E^{\prime}(u),\Pi_{h}u-u\rangle$, we obtain
@@
At this stage, we select $u=u^{*}$ and note that for any $v\in K$, 60 and 9 provide the bound
@@
Thus, adding $\langle E^{\prime}(u^{\ast}),v-u_{h}^{\ell}\rangle$ to both sides of 67 and recalling that $Bu^{*}=o^{*}$, we arrive at
@@
We first estimate $\|B^{\prime}(\lambda^{*}-\lambda_{h}^{\ell})\|_{V^{\prime}}$. For any $v\in V$, we use the Fortin operator 43 and equation 10a to write
@@
Since the map $B:V\rightarrow Q$ is surjective, it follows from the closed range Theorem, see e.g. [24, Lemma A.40], that there exists a constant $\beta>0$ with
@@
Combining 70 with 71 finishes the proof. ∎
@@
(72) $\|Bu^{*}-o_{h}^{\ell}\|_{L^{2}(\Omega_{d})}\leq\|B(u^{*}-u_{h}^{\ell})\|_{L^{2}(\Omega_{d})}+\|(I-P_{h})(Bu_{h}^{\ell}-o_{h}^{\ell})\|_{L^{2}(\Omega_{d})},$
where $P_{h}$ is the $L^{2}(\Omega_{d};\mathbb{R}^{m})$-projection operator onto $W_{h}$, defined by
(73) $(P_{h}o,w_{h})_{\Omega_{d}}=(o,w_{h})_{\Omega_{d}}~\text{ for all }o\in L^{2}(\Omega_{d};\mathbb{R}^{m}),~w_{h}\in W_{h}.$
@@
This is a direct consequence of the triangle inequality and the observation that $P_{h}Bu_{h}^{\ell}=P_{h}o_{h}^{\ell}$, which follows from 10b. ∎
@@ 3.5. Convergence rates and mesh-independence
We now derive abstract error convergence rates from Theorem 3.3 and Lemmas 3.4 and 3.5. For simplicity, we focus on the following spaces $V\subset H^{1}(\Omega)$ and $V\subset H^{1}(\Omega;\mathbb{R}^{n})$. Hereafter, we do not explicitly differentiate between scalar and vector-valued spaces, using the notation for the scalar space while noting that the same results hold, mutatis mutandis, for vector-valued spaces. To derive error rates, we require the construction of a reconstruction operator that allows us to control the last term in (59). In turn, we make the following general assumption, which is critical to the definition of this map. First, we motivate this assumption with a remark.
@@
(74) $u_{h}^{\ell}\in K_{h}=\left\{v_{h}\in V_{h}\mid\int_{K}(v_{h}-\phi)\geq 0\right\}.$
As expected from the analysis of nonconforming discretizations of obstacle problems [6, Section 5.2.1], one then requires an operator $\mathcal{E}_{h}:K_{h}\rightarrow K$ with suitable approximation properties. See Section 4.3 for an explicit construction of the reconstruction operator for $K_{h}$ defined in 74.
Assumption 2 (Reconstruction operator).
There exists a continuous map $\mathcal{E}_{h}:V\rightarrow V$ with the property that
(75a) $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{E}_{h}:K\cup K_{h}\rightarrow K$},
where
(75b) $K_{h}:=\{v_{h}\in V_{h}\mid P_{h}Bv_{h}\in P_{h}(\mathcal{O})\},$
and $P_{h}\colon L^{2}(\Omega_{d};\mathbb{R}^{m})\to W_{h}$ is the $L^{2}$-projector onto $W_{h}$ 73.
The set $K_{h}$ contains functions $u_{h}\in V_{h}$ such that the $L^{2}$-projection of $Bu_{h}\in W_{h}$ matches the $L^{2}$-projection of a constrained observable $o\in\mathcal{O}$. It is constructed to ensure that $u_{h}^{k}\in K_{h}$ for all iterations $k$; see 10b. To derive error estimates, we require the following quasi-interpolation assumption on the stability and approximation of the Fortin and reconstruction operators. The decomposition of the reconstruction operator given in 77 reflects the constructions used in Sections 4 and 5, below.
@@
Assume that the Fortin operator satisfying 43 is stable in the sense that
@@
for all $v\in V$. Further, assume that the reconstruction operator $\mathcal{E}_{h}$ of Assumption 2 is affine linear:
@@
Since $E^{\prime}(u^{*})\in H^{r-1}(\Omega)$, we further bound the last two terms in 59 by $\|E^{\prime}(u^{*})\|_{H^{r-1}(\Omega)}\|\Pi_{h}u^{*}-u^{*}\|_{H^{1-r}(\Omega)}$ and by $\|E^{\prime}(u^{*})\|_{H^{r-1}(\Omega)}\|v-u_{h}^{\ell}\|_{H^{1-r}(\Omega)}$ respectively. Thus, we obtain
@@
To handle the $\|u^{*}-\Pi_{h}u^{*}\|_{H^{1-r}(\Omega)}$ term in 80, we first note that from space interpolation between $L^{2}(\Omega)$ and $H^{1}(\Omega)$ (see, e.g., [9, Chapter 14, Proposition 14.1.5]) and 78a, we obtain that for any $w\in H^{1+s}(\Omega)$,
@@
The estimate on $\|\lambda^{*}-\lambda_{h}^{\ell}\|_{H^{-1}(\Omega)}$ follows from Lemma 3.4. In particular, the first term in 69 is bounded by:
@@ 3.6. Discussion of results
  • First, the PG iterates $u^{\ell}_{h}$ converge to the true solution $u^{\ast}$ as $\ell\to\infty$ and $h\to 0$, even with bounded proximity parameters, $\alpha_{k}\leq\textrm{const.}$ This stands in contrast to penalty [32, Chapter 1.7] and interior point methods [42, Chapter 19] — and even augmented Lagrangian methods in infinite-dimensional spaces [39, 3] — which all require taking a relaxation parameter to a singular limit.
  • Second, the number of iterations is asymptotically independent of the mesh size. More specifically, Corollary 1 shows that a user can guarantee convergence to any desired accuracy by independently selecting the mesh size $h$ and iteration count $\ell$. Asymptotic mesh-independence is a desirable feature of a numerical method [49], often observed in, e.g., Newton’s method [1, 54]. Notably, asymptotic mesh-independence is sometimes witnessed in the primal-dual active set method [36], but for a more restrictive class of problems than considered here [38].
  • Finally, PG can be applied without modification to problems with low-regularity solutions or multipliers. Indeed, Theorem 3.6 provides sufficient conditions on $\Pi_{h}$ and $\mathcal{E}_{h}$ to obtain optimal convergence rates in $h$ depending naturally on the solution $u^{\ast}$ and Fréchet Derivative $E^{\prime}(u^{\ast})$. The following two sections show that these abstract conditions are checkable in practice. It is well-known that penalty methods can also be applied without modification in the low-regularity setting [37]. However, to the best of our knowledge, the existing theory for discretized penalty methods requires $u^{*}\in H^{1+s}(\Omega)$ and $s\in(1/2,1]$ when applied to the obstacle and Signorini problems. See [48, 35] for the obstacle problem and [16, Section 6] for the Signorini problem. Notably, the PG results hold for any $s\in(0,1]$.
@@ 4. Application I: The obstacle problem
and the obstacle $\phi\in H^{1}(\Omega)\cap L^{\infty}(\Omega)$ is fixed. For simplicity, we assume that $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi|_{\partial\Omega}=-\delta\leq 0$} for some non–negative constant $\delta$. It is well-known that there exists a unique solution to 87 [18, Theorem 6.9-1], here denoted by $u^{\ast}$.
@@
Assumption 1 holds for this choice of Legendre function [40, Theorem 4.1]; see also [40, Proposition A.8]. Alternative choices for $\mathcal{R}$ are also admissible, without affecting the proceeding error analysis.